can i get help to solve this and simplify
2 + 22 · 19 + 6
and one more
5m + 4 = 24; 4 is it a solution or not a solution?
thanks

Answer by rajagopalan(174)

Question 1: can i get help to solve this and simplify
2 + 22 · 19 + 6
Step 1 =22*19=418
Step 2=2+418=420
Step 3=420+6=426
Answer=426.
**********
Question 2:
5m + 4 = 24; 4 is it a solution or not a solution?
5m+4=24
5m=24-4=20
m=20/5=4
Answer 4 is OK.